OpenStudy (anonymous):

How to solve the Bernoulli differential equation of y'+2xy=xy^2?

OpenStudy (loser66):

y'= xy^2 -2xy=x(y^2 -2y) dy/(y^2 -2y) = xdx partial fraction for the left hand side to easily take integral then you get the answer